# LIFT: Improving Long Context Understanding of Large Language Models through Long Input Fine-Tuning

### v2

@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023), have revolutionized the field of natural language processing, driving breakthroughs in text generation and significant advancements in tasks like translation, summarization, and conversation. Long sequences, which can span up to millions of tokens, are common in real-world applications, including long books (Kočiskỳ et al., 2018), accounting documentsdocumentsdocuments (Li et al., 2024), high-resolution videos (Wu et al., 2024; Tapaswi et al., 2016), and audio signals (Yang et al., 2024). Extending context windows allows models to capture dependencies across larger text spans and improve coherence, understanding, and accuracy in tasks that require reasoning over extended inputs.
@@
To address these challenges, researchers have developed various techniques to improve the long-context abilities of LLMs. Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Xu et al., 2023) and prompt compression (Jiang et al., 2023) aim to preprocess long inputs within a limited short context window by adaptive retrieval or text compression (El-Kassas et al., 2021). However, the effectiveness of these methods depends on the precision and relevance of the contextual information provided within the context window. It will lead to further hallucinations when noisy, ambiguous, or conflicting information is provided. Long-context adaptation focuses on fine-tuning pretrained LLMs on corpora of long texts to extend their context windows (Chen et al., 2023b; Peng et al., 2023) and is more frequently used in more recent works. However, the adaptation process comes with significant costs in terms of both training data and computational resources. Additionally, with the extended context window, the cost of processing and generating long texts grows quadratically with the input length. Finally, despite the extension, the context windows of these LLMs remain finite, preventing them from generalizing to inputs of infinite length.
To address the above challenges, in this paper, we present a novel framework Long Input Fine-Tuning (LIFT), designed to enhance the long-context capabilities of arbitrary (short-context) models by directly adapting model parameters to the long input. Our approach has the following advantages:
  • Efficient long-input training on the fly. LIFT dynamically adapts to newly introduced long inputs as fresh knowledge by adjusting model parameters, thereby eliminating the need for resource-intensive offline long-context adaptation. To enhance memorization of the long input, we segment it into overlapping segments which can be fitted into a short context window and fine-tune the LLM on batches of the segments. Additionally, we improve long-context comprehension and reasoning through fine-tuning on well-designed auxiliary tasks, further optimizing performance on downstream applications.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG (Lewis et al., 2020) improves the performance of long-context understanding by integrating LLMs with external data sources or memory modules for retrieval (Xu et al., 2023; Jiang et al., 2024; Wang et al., 2024; Jin et al., 2024), thereby avoiding the need to process long inputs directly. Its performance heavily relies on the quality of retrieved content, which must be relevant and concise enough to fit within models’ short context windows. RAG can experience significant performance degradation or hallucination issues when the retrieved context is inaccurate or mismatched.
@@ 3.4 Gated Memory Architecture
To efficiently apply LIFT, we aim to use a parameter-efficient fine-tuning (PEFT) method rather than full-parameter fine-tuning. Existing representative PEFT methods such as LoRA (Hu et al., 2021) and PiSSA (Meng et al., 2024) are not specifically designed for long context tasks. Therefore, we propose a novel Gated Memory adapter working very well in the LIFT framework.
@@
Figure 3: The architecture of Gated Memory. The purple part is the added adapter “gated memory” to fit the out-of-context attention; the green part is the original attention module. During training, only the gated memory part is trained. Other parameters are fixed.
@@ 4.1 Setup
Dataset and metrics. We make a comprehensive evaluation of LIFT on three popular long-context benchmarks, including LooGLE (Li et al., 2023), LongBench (Bai et al., 2023), and Needle-In-A-Haystack (NIAH) (Kamradt, 2023) covering a wide variety of application scenarios.
The evaluation metrics are consistent with those used in the original benchmarks. For LongBench and NIAH, the evaluation metrics are task-specific (Zhang et al., 2020). Since most automatic evaluation metrics are sensitive to semantic expression, output format, and length, we utilize GPT4-0613 (Achiam et al., 2023) for LooGLE as recommended in the paper to judge whether the two answers are semantically the same or not, noted as GPT4_score. It has been proven to exhibit high consistency with human evaluation and can serve as a reliable annotator to a great extent (Suri et al., 2023; Liu et al., 2023; Zheng et al., 2023). The prompts implemented can be found in Appendix A.
Models. For open-source LLMs, we select Llama-3-8B-Instruct (Dubey et al., 2024) and Gemma-2-9B-it (Team, 2024) with 8k context windows. For closed-source commercial LLMs, we choose GPT-3.5-turbo-0125 (Chen et al., 2023a) with a 16k context window. It has shown competitive performance on popular long context benchmarks and can be further fine-tuned through API service. Details of the models and hyper-parameters can be seen in Appendix A.
Settings. In this paper, we mainly evaluate LIFT compared with the truncated ICL performance of the selected LLMs. Specifically: ICL denotes truncating the long input by retaining only the beginning and end of texts within the context window of the original model. LIFT denotes first fine-tuning the LLM using the Gated Memory adapter (except for GPT-3.5 which uses the default API tuning service) with the objectives in Equations (4) and (5) with the above-mentioned truncated ICL. The number of auxiliary tasks used and the method for generating them are specialized for each subtask (detailed in Appendix A).
@@ 4.2.2 Results on LongBench
Figure 4: GPU time with increasing input length for LIFT with PiSSA or Gated Memory and the ICL baseline. The dashed lines represent the fitted curves, showing linear growth for LIFT and quadratic growth for ICL. The red cross indicates the input length at which the ICL baseline runs out of memory.
@@ 4.2.3 Efficiency
Benefiting from our segmentation strategy (Section 3.1), the computational complexity of our method is linear with respect to the input length. To further evaluate the efficiency of our approach compared to ICL (with full input in the context), we measure the time cost of a single Needle-In-A-Haystack (NIAH) task under both methods. In this experiment, the input lengths are controllable and the primary computational cost stems from processing the input context rather than iterative generation. We compare LIFT with the Gated Memory adapter and another adapter PiSSA (Meng et al., 2024) against ICL, plotting GPU time versus input length along with the fitted curves in Figure 4.
First, we observe that LIFT is significantly more memory-efficient than ICL. Notably, ICL runs out of memory when the input length exceeds 90k tokens on our A800 (80G) system. For ICL, the KV cache consumes most of the memory. In contrast, LIFT is capable of handling arbitrarily long inputs. Our segmentation and truncation strategy ensures that LIFT only involves training and inference with short text segments, eliminating the need for extensive caching.
We empirically verify that the time cost of ICL grows quadratically with input length, while our method scales linearly. However, we also observe that the constant factor introduced by adaptation in the computational complexity of LIFT is non-negligible. As a result, LIFT with PiSSA only surpasses ICL in time efficiency when the input length exceeds a certain threshold above 800k tokens, while LIFT with Gated Memory is even slower. We note that although Gated Memory involves fewer trainable parameters than PiSSA, it results in smaller update step sizes and requires more training epochs. A more hardware-friendly implementation is also lacking. As the primary cost of LIFT arises from multi-epoch fine-tuning, we hypothesize that by employing better parallel fine-tuning techniques, the efficiency of LIFT can be significantly improved.
@@ 6 Limitations and Future Work
Limitations of LIFT without ICL. While we often employ truncated contexts to simplify inference on lengthy texts, this approach is proven insufficient for tasks that demand precise information extraction from extended contexts, such as the Needle in a Haystack (NIAH) task (Appendix B).
@@
Appendix A Experiment Details
A.1 Hardware settings
All the experiments, including the main experiments on LooGLE (Section 4.2.1) and LongBench (Section 4.2.2), the efficiency test (Section 4.2.3), and the Needle-in-A-Haystack task (Appendix B), are conducted on a single NVIDIA A800 Tensor Core GPU. We intentionally select this resource-constrained hardware setup, where full-parameter fine-tuning is impractical. This necessitates the use of parameter-efficient fine-tuning (PEFT) methods, which optimize both time and memory efficiency.
Table 8: Resource costs (GPU hours) of the experiments.
Models Methods LooGLE ShortQA LooGLE LongQA LongBench
Llama3 Baseline 3 3 2
PiSSA 20 42 24
Gated Memory 15 33 21
[…]
A.2 Hyperparameter settings
We design a two-stage training paradigm for both the Gated Memory and PiSSA. The two stages differ in the data used in each. In the first stage, the model is trained solely on segmented long inputs (Section 3.1) and optimizes the loss function $\mathcal{L}_{input}$ (Equation 1). In the second stage, auxiliary tasks (Section 3.2) are incorporated into the training dataset, and the model optimizes the loss function $\mathcal{L}$ (Equation 3).
@@ A.2 Hyperparameter settings
A.3 Generating auxiliary tasks
@@
For PassageRetrievalEN in LongBench, we imitate the generation process of the test set — randomly select a passage and prompt the generator to summarize the passage. The auxiliary task is to answer the index of the passage given the generated summary. The prompt for summarizing a passage is as following:
@@
Instruction for LooGLE: You are given a piece of text as the context. You should generate ONLY one question and the corresponding answer according to the context. You should also select one or more sentences directly from the original context as the evidence. The evidences must be EXACTLY SAME ADJACENT sentences retrieved from the context; KEEP the special tokens in the sentences. Please answer in the following format: Question: [question] Answer: [answer] Evidence: [evidence] Please DON’T output quotes when outputting evidences. The following is the piece of text: {Context}
Instruction for Musique: You are given a piece of text as the context. You should generate ONLY one question and the corresponding answer according to the context. You should also select one or more sentences directly from the original context as the evidence. The evidences must be EXACTLY SAME ADJACENT sentences retrieved from the context; KEEP the special tokens in the sentences. Please answer in the following format: Question: [question] Answer: [answer] Evidence: [evidence] Please DON’T output quotes when outputting evidences. The question should focus on the details like names, dates, e.t.c., and the answer should be as brief as possible. The following is the piece of text: {Context}
Instruction for Narrativeqa: You are given a piece of text as the context. You should generate ONLY one question and the corresponding answer according to the context. You should also select one or more sentences directly from the original context as the evidence. The evidences must be EXACTLY SAME ADJACENT sentences retrieved from the context; KEEP the special tokens in the sentences. Please answer in the following format: Question: [question] Answer: [answer] Evidence: [evidence] Please DON’T output quotes when outputting evidences. The question should focus on the details like names, dates, e.t.c. The following is the piece of text: {Context}
A.4 GPT4_score evaluation
@@
The prompt is as following (in the format of the chat template):
System: Given one question, there is a groundtruth and a predict_answer. Please decide whether they are the same or not in semantic. Please only output ’True’ or ’False’ . User: Question: {Question} groundtruth = {Ground-truth answer} predict_answer = {LLM response}
Appendix B Results on Needle-in-a-Haystack (NIAH)

### v3

@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023), have revolutionized the field of natural language processing, driving breakthroughs in text generation and significant advancements in tasks like translation, summarization, and conversation. Long sequences, which can span up to millions of tokens, are common in real-world applications, including long books (Kočiskỳ et al., 2018), accounting documents (Li et al., 2024), high-resolution videos (Wu et al., 2024; Tapaswi et al., 2016), and audio signals (Yang et al., 2024). Extending context windows allows models to capture dependencies across larger text spans and improves coherence, understanding, and accuracy in tasks that require reasoning over extended inputs.
@@
To address these challenges, researchers have developed various techniques to improve the long-context abilities of LLMs. Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Xu et al., 2023) and prompt compression (Jiang et al., 2023) aim to preprocess long inputs within a limited short context window by adaptive retrieval or text compression (El-Kassas et al., 2021). However, the effectiveness of these methods depends on the precision and relevance of the contextual information provided within the context window. It will lead to further hallucinations when noisy, ambiguous, or conflicting information is provided. Long-context adaptation focuses on fine-tuning pretrained LLMs on corpora of long texts to extend their context windows (Chen et al., 2023b; Peng et al., 2023) and is more frequently used in more recent works. However, the adaptation process comes with significant costs in terms of both training data and computational resources. Additionally, with the extended context window, the cost of processing and generating long texts grows quadratically with the input length. Finally, despite the extension, the context windows of these LLMs remain finite, preventing them from generalizing to inputs of unbounded length.
To address the above challenges, in this paper, we present a novel framework, Long Input Fine-Tuning (LIFT), designed to enhance the long-context capabilities of arbitrary (short-context) models by directly adapting model parameters to the long input. Our approach has the following advantages:
  • Efficient long-input training on the fly. LIFT dynamically adapts to newly introduced long inputs as fresh knowledge by adjusting model parameters, thereby eliminating the need for resource-intensive offline long-context adaptation. To enhance memorization of the long input, we segment it into overlapping segments which can be fitted into a short context window and fine-tune the LLM on batches of the segments. Additionally, we improve long-context comprehension and reasoning by fine-tuning on well-designed auxiliary tasks and moving from the naive language modeling task to the contextualized training task, further optimizing performance on downstream applications.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG (Lewis et al., 2020) improves the performance of long-context understanding by integrating LLMs with external data sources or memory modules for retrieval (Xu et al., 2023; Jiang et al., 2024; Wang et al., 2024a; Jin et al., 2024), thereby avoiding the need to process long inputs directly. Its performance heavily relies on the quality of retrieved content, which must be relevant and concise enough to fit within models’ short context windows. RAG can experience significant performance degradation or hallucination issues when the retrieved context is inaccurate or mismatched.
@@
Wang et al. (2024b) explore how TTT can enhance LLMs in long generation tasks such as novel writing and translation through iteratively fine-tuning a LoRA adapter to memorize the previously generated tokens. While sharing a similar idea to store context knowledge in LLM parameters, LIFT focuses on long-context understanding which poses different challenges, such as accurate memorization of details and capturing intricate long/short dependencies. Another important difference is that LIFT adopts a novel Gated Memory Adapter instead of LoRA. Traditional adapters like LoRA modifies model weights by directly adding delta weights, which unanimously changes the outputs for all inputs. In contrast, our Gated Memory Adapter uses a gate to control how much to use the original model and the adapter, dynamically determining whether to use pre-trained knowledge or new knowledge. Experiments confirm its great advantages.
@@ 3.4 Gated Memory Architecture
To efficiently apply LIFT, we aim to use a parameter-efficient fine-tuning (PEFT) method rather than full-parameter fine-tuning. Existing representative PEFT methods such as LoRA (Hu et al., 2021) and PiSSA (Meng et al., 2024), as a better parameter initialization for LoRA, are not specifically designed for long context tasks. Therefore, we propose a novel Gated Memory adapter working very well in the LIFT framework.
Figure 3: The architecture of Gated Memory. The purple part is the added adapter “gated memory” to fit the out-of-context attention; the green part is the original attention module. During training, only the gated memory part is trained.
@@ 4.1 Setup
Dataset and metrics. We evaluate LIFT on two popular long-context benchmarks, including LooGLE (Li et al., 2023) and LongBench (Bai et al., 2023), covering a wide variety of application scenarios.
The evaluation metrics are consistent with those used in the original benchmarks. For LongBench and NIAH, the evaluation metrics are task-specific (Zhang et al., 2020). Since most automatic evaluation metrics are sensitive to semantic expression, output format, and length, we utilize GPT4-0613 (Achiam et al., 2023) for LooGLE as recommended in the paper to judge whether the two answers are semantically the same or not, noted as GPT4_score. It has been proven to exhibit high consistency with human evaluation and can serve as a reliable annotator to a great extent (Suri et al., 2023; Liu et al., 2023; Zheng et al., 2023). The prompts implemented can be found in Section B.2.
Models. For open-source LLMs, we select Llama-3-8B-Instruct, Llama-3.1-8B-Instruct (Dubey et al., 2024), and Gemma-2-9B-it (Team, 2024) with 8k context windows. For closed-source commercial LLMs, we choose GPT-3.5-turbo-0125 (Chen et al., 2023a) with a 16k context window. It has shown competitive performance on popular long context benchmarks and can be fine-tuned through API service.
Settings. In this paper, we mainly evaluate LIFT compared with the truncated ICL performance of the selected LLMs. Specifically: ICL denotes truncating the long input by retaining only the beginning and end of texts within the context window of the original model. LIFT denotes first fine-tuning the LLM using the Gated Memory adapter (except for GPT-3.5 which uses the default API tuning service) with the objectives in Equations (4) and (5) with the above-mentioned truncated ICL.
For efficiency comparison and the ablation study on Gated Memory, we conduct a comparison between Gated Memory and PiSSA (Meng et al.). Specifically, using the same training framework (including task designs, objective functions, and scheduling configurations) of LIFT, we fine-tune the model separately with the Gated Memory adapter and the PiSSA adapter. Additionally, we also evaluate the efficiency under an ideal setting, denoted as LIFT (Llama 3.1), where we apply LIFT to Llama-3.1-8B-Instruct and assume that all the knowledge in the inputs is absorbed into the parameters and truncated ICL is no longer needed.
The implementation details can be seen in Appendix A.
@@ 4.2.2 Results on LongBench
Figure 4: Subfigures (a)-(c) illustrate the decoding speed comparison between LIFT and ICL given inputs of length 20K, 50K, and 100K.
@@ 4.2.3 Efficiency
Benefiting from transferring input tokens into LLM parameters, LIFT alleviates the need to compute the attention score with all the input tokens when generating a token. Consequently, the overall efficiency of LIFT is much higher than that of long-context ICL. To further evaluate the efficiency of our approach compared to ICL (with full input in the context), we measure the time cost of generating $y$ tokens with $x$ tokens as the input. Empirically, as illustrated in Figure 4 (a)-(c), LIFT starts to outperform ICL in decoding speed when generating more than 1500 tokens. This is because, LIFT only needs to fine-tune on the long input once, and later fully becomes a short-context model with very short decoding time per token. In contrast, ICL puts all the long input in the context, and every new token generation needs to compute the attention of the last token to all the input tokens, incurring great latency in every new token generation.
@@ 6 Limitations and Future Work
Limitations of LIFT without ICL. While we often employ truncated contexts to simplify inference on lengthy texts, this approach is proven insufficient for tasks that demand precise information extraction from extended contexts, such as the Needle in a Haystack (NIAH) task (Appendix C).
@@
Appendix A Implementation Details
A.1 Training data
The training data of LIFT includes contextualized input segments (Sections 3.1 and 3.3) and auxiliary tasks (Section 3.2). For contextualized input segments, the length of each input segment is $\ell=2048$, and the offset between two segments is $s=\frac{3}{8}\ell=768$ (Figure 2). We affiliate each segment with an extra context to simulate ICL during testing. We randomly select $r\in[0,4096]$ ($r$ is sampled independently for each segment) and concatenate the first $r$ tokens and the last $(4096-r)$ tokens of the long input as the context. The prompt is as follows:
{First $r$ tokens}…{Last $(4096-r)$ tokens}Given above context, please recite following segment of the context: {The segment}
Only {The segment} is supervised during fine-tuning.
For each subtask of LooGLE and LongBench, except PassageRetrievalEN, to generate auxiliary tasks, we randomly select 16 consecutive sentences from the long input as the context and prompt LLM to synthesize a question and the corresponding answer based on the context. Details of the prompt used can be found in Appendix B.
A.2 Training process and hyperparameters
We design a two-stage training paradigm for both Gated Memory and PiSSA. In the first stage, the model is trained solely on contextualized input segments and optimizes the loss function $\mathcal{L}_{input}$ (Equation 1). In the second stage, auxiliary tasks (Section 3.2) are incorporated with contextualized input segments as the training data, and the model optimizes the loss function $\mathcal{L}$ (Equation 3).
@@ A.2 Training process and hyperparameters
A.3 Hardwares
All the experiments, including the main experiments on LooGLE (Section 4.2.1) and LongBench (Section 4.2.2), the efficiency test (Section 4.2.3), and the Needle-in-A-Haystack task (Appendix C), are conducted on a single NVIDIA A800 Tensor Core GPU. We intentionally select this resource-constrained hardware setup, where full-parameter fine-tuning is impractical. This necessitates the use of parameter-efficient fine-tuning (PEFT) methods, which optimize both time and memory efficiency.
Table 10: Resource costs (GPU hours) of the experiments.
Models Methods LooGLE ShortQA LooGLE LongQA LongBench
Llama3 ICL 3 3 2
PiSSA 20 42 24
Gated Memory 15 33 21
[…]
Appendix B Prompts
B.1 Auxiliary tasks generation
@@
Instruction for LooGLE:
You are given a piece of text as the context. You should generate ONLY one question and the corresponding answer according to the context. You should also select one or more sentences directly from the original context as the evidence. The evidences must be EXACTLY SAME ADJACENT sentences retrieved from the context; KEEP the special tokens in the sentences. Please answer in the following format:
Question: [question] Answer: [answer] Evidence: [evidence] Please DON’T output quotes when outputting evidences. The following is the piece of text: {Context}
Instruction for Musique:
You are given a piece of text as the context. You should generate ONLY one question and the corresponding answer according to the context. You should also select one or more sentences directly from the original context as the evidence. The evidences must be EXACTLY SAME ADJACENT sentences retrieved from the context; KEEP the special tokens in the sentences. Please answer in the following format:
Question: [question] Answer: [answer] Evidence: [evidence] Please DON’T output quotes when outputting evidences. The question should focus on the details like names, dates, e.t.c., and the answer should be as brief as possible. The following is the piece of text: {Context}
Instruction for Narrativeqa:
You are given a piece of text as the context. You should generate ONLY one question and the corresponding answer according to the context. You should also select one or more sentences directly from the original context as the evidence. The evidences must be EXACTLY SAME ADJACENT sentences retrieved from the context; KEEP the special tokens in the sentences. Please answer in the following format:
Question: [question] Answer: [answer] Evidence: [evidence] Please DON’T output quotes when outputting evidences. The question should focus on the details like names, dates, e.t.c. The following is the piece of text: {Context}
For PassageRetrievalEN in LongBench, we imitate the generation process of the test set by randomly selecting a passage and summarize the passage by LLM. The auxiliary task is to answer the index of the passage given the generated summary.
Instruction for passage summarization:
@@
B.2 GPT4_score evaluation
@@
Given one question, there is a groundtruth and a predict_answer. Please decide whether they are the same or not in semantic. Please only output ’True’ or ’False’ . Question: {Question} groundtruth = {Ground-truth answer} predict_answer = {LLM response}
Appendix C Results on Needle-in-a-Haystack (NIAH)
@@
Appendix D Analysis to the end-to-end training of Gated Memory
To train Gated Memory, we learn two MLPs, $g(q)$ and $m(q)$. Suppose we have a complete input $(\mathbf{x}^{\prime},\mathbf{x})$, where $\mathbf{x}^{\prime}$ denotes the full long input (e.g., a long document) and $\mathbf{x}$ denotes the prompt (e.g., a question). Let the hidden states of the last layer be $(\hat{\mathbf{h}}^{\prime(l-1)},\hat{\mathbf{h}}^{(l)})$, corresponding to $(\mathbf{x}^{\prime},\mathbf{x})$, and let the queries, keys, and values be $\hat{\mathbf{q}},\hat{\mathbf{k}},\hat{\mathbf{v}}$. Equation 7 defines the ground truth for $g(q_{L})$ and $m(q_{L})$, where $l$ and $l^{\prime}$ are the length of $\mathbf{x}$ and $\mathbf{x}^{\prime}$, respectively. However, directly using Equation 7 as supervision (distillation) is infeasible as passing the entire long input $\mathbf{x}^{\prime}$ through the model is too expensive. Instead, we treat Gated Memory as an adapter similar to LoRA and train it in a SFT manner (end-to-end training).
Empirically, the end-to-end training proves effective, outperforming ICL and PiSSA on most tasks (Tables 2, 3 and 6). However, it doesn’t fully replicate the theoretical attention mechanism. We adopt Llama-3.1-8B-Instruct with a 128K context window and compare the attention outputs under the following settings:
  • Ground-truth. The model has access to the complete context.
  • Base model. The model can only see the question.
  • Gated Memory. The model is fine-tuned using the LIFT training process with the Gated Memory adapter. During evaluation, the model can only see the question.
In detail, the input text for each setting is as follows:
Ground-truth:
Given the article ”{Title}”: {Context} Based on the article {Title}, please answer the following question: {Question}
Answer: {Answer} Base model and Gated Memory: Based on the article {Title}, please answer the following question: {Question} Answer: {Answer}
In the experiment, {Title} and {Context} are the title and the content of the first document in LooGLE ShortQA dataset, and {Question} and {Answer} are the questions and ground-truth answers provided by LooGLE. We collect only the attention outputs of the {Answer} tokens and compute the cosine similarity between the attention output of Ground-truth and those of Base model and Gated Memory, respectively, as illustrated in Figure 6. The similarity corresponding to Gated Memory is consistently lower than that corresponding to Base model, suggesting that the end-to-end trained Gated Memory learns to answer questions through mechanisms other than full-context ICL. We consider conducting more comprehensive experiments with full-attention distillation as an important direction for future work.
Figure 6: The cosine similarity between the Ground-truth attention output and the Gated Memory and Base model output respectively.